OpenStudy (anonymous):

Solve by using quadratic formula r^2+2r=1

hartnn (hartnn):

subtract 1 from both sides to get r^2+2r-1=0 Compare your quadratic equation with \(ar^2+br+c=0\) find \[a=...?\\b=...?\\c=...?\] then the two roots of r are: \(\huge{r_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

a=1 b=2 c=-1 right?

hartnn (hartnn):

correct :) now just plug it in the formula

OpenStudy (anonymous):

Hmm i can't type the equation where i'm stuck

OpenStudy (anonymous):

I'm only using mobile, it's hard to type the equation

hartnn (hartnn):

lets go step by step, \(b^2-4ac=...?\) just give final values..

OpenStudy (anonymous):

8

hartnn (hartnn):

good, so numerator = \(-2 \pm \sqrt 8\) right ? can you simplify \(\sqrt 8\) ??

OpenStudy (anonymous):

2 square root of 2

OpenStudy (anonymous):

I can't use symbols sorry

hartnn (hartnn):

i understood that , its \(2 \sqrt2\)and is correct :) so, \(\huge{x_{1,2}=\frac{-2 \pm 2\sqrt2}{2} = -1 \pm \sqrt2}\) so the 2 roots are -1+ root 2, and -1- root 2

OpenStudy (anonymous):

I can't see the whole equation

hartnn (hartnn):

r1,2 = (-2 +or - 2sqrt2 )/2=-1 + or - sqrt 2 so, 2 roots are r1 = -1+sqrt 2 r2 =-1-sqrt2

OpenStudy (anonymous):

How did it become -1?

hartnn (hartnn):

factoring out 2 in the numerator (-2 +or - 2sqrt2 )/2 = 2 (-1 +or - sqrt 2)/2 the 2 in the numerator and denominator cancels out

OpenStudy (anonymous):

Oh i see, sorey, i get it :£

OpenStudy (anonymous):

:)

hartnn (hartnn):

great ! :)